9. Prove that (a - b) x (a+ b) = 2(a x b) X
Answer #1

I am attaching the pictures of my solution. If you find my answer helpful please press thumbs up button.

A To proeve a- x (a) = 9(TXB) a a,1 let 4,3ask -b bk + a-E la,i+ a, 3 +a,k)- (b,i +b,5+by) -B = (a,-b) 1 +(a,-b.)+-)k (a,+b ((aB)xB)-aa +a b b4-b, b-4a,-b,a, tb4, tb,by So +3(as4, +4,b, -4,b -,b-4,4,-4, by +b, 03 tb,bs R4,43a b-b,a -b,a,4,-b.4 +abtbb